# Invariant Subspaces of Perturbed Backward Shift

### v1

@@ 2. Perturbations of $T_{\varphi}^{*}$
for all $z\in\mathbb{T}$ a.e. (in the sense of radial limits of $H^{\infty}$-functions). The analytic Teoplitz operator with symbol $\varphi\in H^{\infty}$ is then defined by
@@
for all $f\in H^{2}$. We will consider rank-$m$ perturbations of the Toepliz operator $T_{\varphi}$. Let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be arbitrary but fixed orthonormal sets of vectors in $H^{2}$. Define the rank-$m$ perturbation $\mathcal{T}_{[\varphi;u,v]}$ of $T_{\varphi}^{*}$ on $H^{2}$ by
@@
Returning to our perturbation situation, we first define Sarason-type finite-rank perturbation $\mathcal{T}_{[\varphi,u]}$ on $H^{2}$ as
@@
where $\mathcal{F}=span\{u_{1},u_{2},\ldots,u_{m}\}$ is a finite-dimensional subspace of $H^{2}$. Since $T_{\varphi}$ is a contraction and $I-P_{\mathcal{F}}$ is an orthogonal projection, the operator $\mathcal{T}_{[\varphi,u]}$ is a contraction. Since $T_{\varphi}\in C_{\cdot 0}$ and $\mathcal{F}$ is finite-dimensional, (2.1) stated above implies that $T_{\varphi}(I-P_{\mathcal{F}})$ is a $C_{\cdot 0}$ contraction. For each $g\in H^{2}$ and $n\geq 1$, we compute
@@
classifies $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$-invariant closed subspaces of $H^{2}_{\mathbb{C}^{p}}$. This is the consequence of the celebrated Beurling-Lax theorem. The subspace $\mathcal{K}_{\Psi}$ is commonly referred to as the model space corresponding to the inner function $\Psi$.
Now we are in a position to present the description of invariant subspaces of $T_{\varphi}^{*}$. In the following statement, $[f_{1},\ldots,f_{p},1]_{1\times(p+1)}$ refers to the row vector in the $(p+1)$-copies of $H^{2}$ (which is identified with $H^{2}_{\mathbb{C}^{p+1}}$). Moreover, often we will express vectors $F$ in $H^{2}_{\mathbb{C}^{p}}$ as column vectors $[f_{1},\ldots,f_{p}]^{t}$, where $f_{i}\in H^{2}$ for all $i=1,\ldots,p$.
@@
$\mathcal{K}:=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|(F,f_{0})\|^{2}\}.$
@@
We open by noting that the following inequality occurs within the context of the theorem’s statement:
$p:=\dim\mathcal{W}\leq m=\text{rank }(\sum_{i=1}^{m}v_{i}\otimes u_{i}),$
where following our convention, we simply write $\mathcal{T}_{[\varphi;u,v]}=T_{\varphi}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, and define the subspace $\mathcal{W}:=span\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Assume that $\mathcal{M}$ is invariant under $\mathcal{T}_{[\varphi;u,v]}$. If $\mathcal{W}=\{0\}$, then $\{u_{1},u_{2},\ldots,u_{m}\}\perp\mathcal{M}$, which implies that $\mathcal{M}$ is $T_{\varphi}^{*}$ invariant. In this case, we simply write $\mathcal{M}=[1]\mathcal{K}$, where $\mathcal{K}=\mathcal{M}$. Next, we assume that
@@
Let
$dim(\mathcal{W})=p.$
Then $1\leq p\leq m$. Pick an orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, and denote the row operator
$F_{0}=[f_{1},f_{2},\ldots,f_{p}]_{1\times p}.$
Decompose $\mathcal{M}$ as
@@
Fix $f\in\mathcal{M}$ and write
$f=P_{\mathcal{W}}(f)\oplus P_{\mathcal{M}\cap\mathcal{W}^{\perp}}(f)=(a_{01}f_{1}+\cdots+a_{0p}f_{p})\oplus g_{1},$
that is
$f=F_{0}A_{0}\oplus g_{1},$ (2.5)
@@
$l_{1}=(T_{\varphi}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i})g_{1}=T_{\varphi}^{*}g_{1}-\sum_{i=1}^{m}\langle g_{1},u_{i}\rangle v_{i}=T_{\varphi}^{*}g_{1},$
as $g_{1}\in\mathcal{M}\cap\mathcal{W}^{\perp}$ and $\mathcal{W}=\text{span}\{P_{\mathcal{M}}u_{j}:j=1,\ldots,m\}$. Since $T_{\varphi}T_{\varphi}^{*}=I-P_{\mathcal{K}_{\varphi}}$, we conclude that $T_{\varphi}l_{1}=T_{\varphi}T_{\varphi}^{*}g_{1}=g_{1}-P_{\mathcal{K}_{\varphi}}g_{1}$. This readily implies that
$g_{1}=\varphi l_{1}+P_{\mathcal{K}_{\varphi}}g_{1},$
@@
$\begin{cases}f=F_{0}A_{0}+l_{1}\varphi+P_{\mathcal{K}_{\varphi}}g_{1}\\ \|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{l_{1}}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{1}\|^{2},\end{cases}$
@@
where $g_{2}=P_{\mathcal{M}\cap\mathcal{W}^{\perp}}l_{1}\in\mathcal{M}\cap\mathcal{W}^{\perp}$, $A_{1}\in\mathbb{C}^{p}$, and
@@
$g_{2}=l_{2}\varphi+P_{\mathcal{K}_{\varphi}}g_{2},$
@@
and $\{A_{i}\}_{i=1}^{n}\subseteq\mathbb{C}^{r}$ for all $n\geq 1$. However, since $\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}=\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}$, it follows that
@@
$\|{l_{n+1}}\|=\|{(\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})^{m+1}f}\|\leq\|{\mathcal{T}_{[\varphi,u]}}\|\|{(P_{\mathcal{M}\cap\mathcal{W}^{\perp}}\mathcal{T}_{[\varphi,u]})^{m}f}\|,$
@@
$\max\Big{\{}\sum_{i=0}^{\infty}\|{A_{i}}\|^{2},\sum_{j=1}^{\infty}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}\Big{\}}\leq\|{f}\|^{2},$
@@
for all $z\in\mathbb{D}$. We claim that $F\in H^{2}_{\mathbb{C}^{p}}$ and $f_{0}\in H^{2}$. Observe that $\ker T_{\varphi}^{*}=\mathcal{K}_{\varphi}$. Because $T_{\varphi}$ is a pure isometry (that is, a shift operator), by the von Neumann-Wold decomposition theorem, we have the orthogonal decomposition
@@
Since $\sum_{j=0}^{\infty}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}<\infty$, in view of the above decomposition, we conclude that $f_{0}\in H^{2}$. For $F$, we set
@@
where $\mathcal{K}_{\Phi}=\ker T_{\Phi}^{*}$. Again, since $T_{\Phi}$ on $H^{2}\otimes\mathbb{C}^{p}$ is a pure isometry (as $T_{\Phi}^{*m}=T^{*m}_{\varphi}\otimes I_{\mathbb{C}^{p}}\rightarrow 0$ in SOT), we again have the von Neumann-Wold decomposition as
@@
in $H^{1}(\mathbb{D})$. Also, we compute
@@
This along with $\|{f-F_{0}K_{n}-k_{n}}\|_{1}\rightarrow 0$ implies that $f=F_{0}F+f_{0}$ in $H^{1}(\mathbb{D})$. But $f\in H^{2}$, and then, finally we get the desired representation of $f\in\mathcal{M}$ as
@@
Therefore, $\tilde{F}=\hat{F}$ and $\tilde{f}_{0}=\hat{f}_{0}$, and hence, under the norm condition (2.8), the representations of elements of $\mathcal{M}$ as in (2.7) are unique. In view of this, now we define the closed subspace $\mathcal{K}$ of $H^{2}_{\mathbb{C}^{p}}\oplus H^{2}\cong H^{2}_{\mathbb{C}^{p+1}}$ as follows:
$\mathcal{K}:=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|{f_{0}}\|^{2}=\|(F,f_{0})\|^{2}\}.$
We claim that $\mathcal{K}$ is invariant under $(T_{\varphi}\otimes I_{\mathbb{C}^{p+1}})^{*}$, that is, $(T_{\varphi}^{*}\otimes I_{\mathbb{C}^{p+1}})\mathcal{K}\subseteq\mathcal{K}$. To see this, pick $(F,f_{1})\in\mathcal{K}$. By the construction of $\mathcal{K}$, there exists $g\in\mathcal{M}$ such that
@@
By the uniqueness part of $g$ as in (2.7) under the condition (2.8), it follows that $F(z)=\sum_{i=0}^{\infty}A_{i}\varphi^{i}$, and $f_{1}(z)=\sum_{j=0}^{\infty}(P_{\mathcal{K}_{\varphi}}g_{j})\varphi^{j}$. In the following, we shall identify $H^{2}$ functions with their radial limit representations. From this point of view, it follows that $\varphi\bar{\varphi}=1$, and hence
@@
$=F_{0}K_{0}(0)+F_{0}\varphi\bar{\varphi}\left(F-F(0)\right)+\varphi\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right)+P_{\mathcal{K}_{\varphi}}f_{0}$
@@
From the representations of $F$ and $f_{0}$, it is easy to see that $T_{\Phi}^{*}F=\bar{\varphi}(F-F_{0}(0))$ and $T_{\varphi}^{*}f_{0}=\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right)$, and hence
$f=FF(0)+\varphi(FT_{\Phi}^{*}F+T_{\varphi}^{*}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}.$
@@
$f=FA_{0}+\varphi(F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}.$
@@
Comparing this with the above representation of $f$, we obtain (recall that $\varphi\bar{\varphi}=1$) $F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0}=l_{1}\in\mathcal{M}$. In other words
@@
Now we apply Theorem 2.1 to the case $\theta(z)=z$ for all $z\in\mathbb{D}$. We already pointed out that the $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$-invariant subspaces are precisely model spaces (see (2.4)). Therefore, we have the following:
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
@@
where $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$

### v2

@@ 2. Perturbations of $T_{\varphi}^{*}$
for a.e. $z\in\mathbb{T}$ (in the sense of radial limits of $H^{\infty}$-functions). Let $\varphi\in H^{\infty}$ be an inner function. In this case, the Toeplitz operator $T_{\varphi}$ (see (1.6)) turns into an analytic Toeplitz operator with the symbol $\varphi\in H^{\infty}$, therefore simplifying to
@@
for all $f\in H^{2}$. We will consider rank-$m$ perturbations of $T_{\varphi}^{*}$. Let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be arbitrary but fixed orthonormal sets of functions in $H^{2}$. Define
@@
Returning to our perturbation situation, we first define Sarason-type finite-rank perturbation $\mathcal{T}_{[\varphi,u]}$ on $H^{2}$ as (see (1.7))
@@
where $\mathcal{F}=span\{u_{1},\ldots,u_{m}\}$ is a finite-dimensional subspace of $H^{2}$. As $T_{\varphi}$ is a contraction and $I-P_{\mathcal{F}}$ is an orthogonal projection, $\mathcal{T}_{[\varphi,u]}$ is a contraction. Since $T_{\varphi}\in C_{\cdot 0}$ and $\mathcal{F}$ is finite-dimensional, (2.1) stated above implies that $T_{\varphi}(I-P_{\mathcal{F}})$ is a $C_{\cdot 0}$ contraction. For each $g\in H^{2}$ and $n\geq 1$, we compute
@@
classifies $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$-invariant closed subspaces of $H^{2}_{\mathbb{C}^{p}}$. This is the consequence of the celebrated Beurling-Lax theorem [14, Chapter V. Theorem 3.3]. The subspace $\mathcal{K}_{\Psi}$ is commonly referred to as the model space corresponding to the inner function $\Psi$ [16].
Now we are in a position to present the description of the invariant subspaces of finite-rank perturbations of $T_{\varphi}^{*}$. In the following statement, $[f_{1},\ldots,f_{p},1]_{1\times(p+1)}$ refers to the row vector in the $(p+1)$-copies of $H^{2}$ (which is identified with $H^{2}_{\mathbb{C}^{p+1}}$). Moreover, often we will express vectors $G$ in $H^{2}_{\mathbb{C}^{p+1}}$ as column vectors $[g_{1},\ldots,g_{p+1}]^{t}$, where $g_{i}\in H^{2}$ for all $i=1,\ldots,p+1$. With this notation in mind, we get
$[f_{1},\ldots,f_{p},1]_{1\times(p+1)}G=\sum_{i=1}^{p}f_{i}g_{i}+g_{p+1},$
whenever the sum is defined.
@@
$\mathcal{K}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$
@@
We open by defining the subspace
$\mathcal{W}:=span\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\},$
and noting that the following inequality occurs within the context of the theorem’s statement:
$p:=\dim\mathcal{W}\leq m=\text{rank }(\sum_{i=1}^{m}v_{i}\otimes u_{i}).$
Following our convention, we set $\mathcal{T}_{[\varphi;u,v]}=T_{\varphi}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$ and $\mathcal{T}_{[\varphi,u]}=T_{\varphi}^{*}-\sum_{i=1}^{m}T_{\varphi}^{*}u_{i}\otimes u_{i}$. We know that $\mathcal{M}$ is invariant under $\mathcal{T}_{[\varphi;u,v]}$. If $\mathcal{W}=\{0\}$, then $\{u_{1},\ldots,u_{m}\}\perp\mathcal{M}$, which implies that $\mathcal{M}$ is $T_{\varphi}^{*}$ invariant. In this case, we simply write $\mathcal{M}=[1]\mathcal{K}$, where $\mathcal{K}=\mathcal{M}$. Next, we assume that $\mathcal{W}$ is nontrivial:
@@
Then $1\leq p\leq m$. By noting that $\mathcal{W}\subseteq\mathcal{M}$, we decompose $\mathcal{M}$ as
@@
Pick an orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, and fix $f\in\mathcal{M}$. There exists $\{a_{0j}\}_{j=1}^{p}\subseteq\mathbb{C}$ such that
$P_{\mathcal{W}}f=a_{01}f_{1}+\cdots+a_{0p}f_{p}.$
We write
$P_{\mathcal{W}}f=F_{0}A_{0},$
@@
$F_{0}=[f_{1},\ldots,f_{p}]_{1\times p},$
denotes the row vector. Since $f=P_{\mathcal{W}}f\oplus P_{\mathcal{M}\cap\mathcal{W}^{\perp}}f$, it follows that
$f=F_{0}A_{0}\oplus g_{1},$ (2.5)
where
@@
$\begin{split}l_{1}&=(T_{\varphi}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i})g_{1}\\ &=T_{\varphi}^{*}g_{1}-\sum_{i=1}^{m}\langle g_{1},u_{i}\rangle v_{i}\\ &=T_{\varphi}^{*}g_{1},\end{split}$
as $g_{1}\in\mathcal{M}\cap\mathcal{W}^{\perp}$ and $\mathcal{W}=\text{span}\{P_{\mathcal{M}}u_{j}:j=1,\ldots,m\}$. Since $T_{\varphi}T_{\varphi}^{*}=I-P_{\mathcal{K}_{\varphi}}$, we conclude that $T_{\varphi}l_{1}=T_{\varphi}T_{\varphi}^{*}g_{1}=g_{1}-P_{\mathcal{K}_{\varphi}}g_{1}$. This readily implies that (note that $T_{\varphi}l_{1}=\varphi l_{1}$)
$g_{1}=\varphi l_{1}\oplus P_{\mathcal{K}_{\varphi}}g_{1},$
@@
$\begin{cases}f=F_{0}A_{0}\oplus l_{1}\varphi\oplus P_{\mathcal{K}_{\varphi}}g_{1}\\ \|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{l_{1}}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{1}\|^{2},\end{cases}$
@@
where $g_{2}=P_{\mathcal{M}\cap\mathcal{W}^{\perp}}l_{1}\in\mathcal{M}\cap\mathcal{W}^{\perp}$, $A_{1}=[a_{11},\ldots,a_{1p}]^{t}\in\mathbb{C}^{p}$, and
@@
$g_{2}=l_{2}\varphi\oplus P_{\mathcal{K}_{\varphi}}g_{2},$
@@
and $\{A_{i}\}_{i=1}^{n}\subseteq\mathbb{C}^{p}$ for all $n\geq 1$. On the other hand, $P_{\mathcal{M}\cap\mathcal{W}^{\perp}}=P_{\mathcal{M}}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}$ and $P_{\mathcal{M}}u_{i}\in\mathcal{W}$ implies that
$(v_{i}\otimes u_{i})P_{\mathcal{M}\cap\mathcal{W}^{\perp}}=0=(T_{\varphi}^{*}u_{i}\otimes u_{i})P_{\mathcal{M}\cap\mathcal{W}^{\perp}},$
for all $i=1,\ldots,m$, and hence
$\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}=\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}(=T_{\varphi}^{*}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}).$
It follows that
@@
$\|{l_{n+1}}\|=\|{(\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})^{n+1}f}\|\leq\|{\mathcal{T}_{[\varphi,u]}}\|\|{(P_{\mathcal{M}\cap\mathcal{W}^{\perp}}\mathcal{T}_{[\varphi,u]})^{n}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}f}\|,$
@@
$\max\Big{\{}\sum_{i=0}^{\infty}\|{A_{i}}\|^{2},\sum_{j=1}^{\infty}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}\Big{\}}\leq\|{f}\|^{2}.$
@@
for all $z\in\mathbb{D}$. We claim that $F\in H^{2}_{\mathbb{C}^{p}}$ and $f_{0}\in H^{2}$. Observe that $\ker T_{\varphi}^{*}=\mathcal{K}_{\varphi}$. Because $T_{\varphi}$ is a pure isometry (that is, a shift operator), by the von Neumann-Wold decomposition theorem [14, Chapter I. Theorem 1.1], we have the orthogonal decomposition
@@
Since $\sum_{j=0}^{\infty}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}<\infty$, in view of the above decomposition, we conclude that $f_{0}\in H^{2}$. For the $\mathbb{C}^{p}$-valued function $F$, we set
@@
where $\mathcal{K}_{\Phi}=\ker T_{\Phi}^{*}$. Again, since $T_{\Phi}$ on $H^{2}\otimes\mathbb{C}^{p}$ is a pure isometry (as $T_{\Phi}^{*n}=T^{*n}_{\varphi}\otimes I_{\mathbb{C}^{p}}\rightarrow 0$ in SOT), we have the von Neumann-Wold decomposition as
@@
in $H^{1}$. Recall that $H^{1}$ is the Banach space of analytic functions on $\mathbb{D}$ defined by [16, Volume 1. Chapter 3]
$H^{1}=\{f\in\text{Hol}(\mathbb{D}):\|f\|:=\sup_{0\leq r<1}\int_{0}^{2\pi}|f(re^{i\theta})|\frac{d\theta}{2\pi}<\infty\}.$
Also, we compute
@@
This along with $\|{f-F_{0}K_{n}-k_{n}}\|_{1}\rightarrow 0$ implies that $f=F_{0}F+f_{0}$ in $H^{1}$. But $f\in H^{2}$, and then, we finally get the desired representation of $f\in\mathcal{M}$ as
@@
Therefore, $\tilde{F}=\hat{F}$ and $\tilde{f}_{0}=\hat{f}_{0}$, and hence, under the norm condition (2.8), the representations of elements of $\mathcal{M}$ as in (2.7) are unique. In view of this, now we define the closed subspace $\mathcal{K}$ of $H^{2}_{\mathbb{C}^{p}}\oplus H^{2}\cong H^{2}_{\mathbb{C}^{p+1}}$ as
$\mathcal{K}:=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|{f_{0}}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$
We claim that $\mathcal{K}$ is invariant under $(T_{\varphi}\otimes I_{\mathbb{C}^{p+1}})^{*}$. To see this, pick $(F,f_{0})\in\mathcal{K}$. By the construction of $\mathcal{K}$, there exists $g\in\mathcal{M}$ such that
@@
By the uniqueness part of $g$ as in (2.7) under the condition (2.8), it follows that $F(z)=\sum_{i=0}^{\infty}A_{i}\varphi^{i}$, and $f_{0}(z)=\sum_{j=1}^{\infty}(P_{\mathcal{K}_{\varphi}}g_{j})\varphi^{j-1}$. In the following, we shall identify $H^{2}$ functions with their radial limit representations. From this point of view, it follows that
$\varphi\bar{\varphi}=1,$
and hence
@@
$=F_{0}F(0)+F_{0}\varphi\bar{\varphi}\left(F-F(0)\right)+\varphi\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right)+P_{\mathcal{K}_{\varphi}}f_{0}$
@@
From the representations of $F$ and $f_{0}$, it is easy to see that $T_{\Phi}^{*}F=\bar{\varphi}(F-F(0))$ and $T_{\varphi}^{*}f_{0}=\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right)$, and hence
$f=F_{0}F(0)+\varphi(F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}.$
@@
$f=F_{0}A_{0}+\varphi(F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}.$
@@
Comparing this with the above representation of $f$, we finally obtain (recall that $\varphi\bar{\varphi}=1$) $F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0}=l_{1}\in\mathcal{M}$. In other words
@@
As we previously mentioned, complete description of the lattices of invariant subspaces of natural operators is a frequently difficult task and a rare occurrence. The answers often offer useful information on a variety of subjects (cf. the Beurling theorem [1] and its applications [14, 16]). At this moment, it is unclear how far the above result extends in terms of application. However, we will continue our discussion in light of the above results. First, we apply it to the case $\theta(z)=z$ for all $z\in\mathbb{D}$. We already pointed out that the $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$-invariant subspaces are precisely model spaces (see (2.4)). Therefore, we have:
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
where $F_{0}=[f_{1},\ldots,f_{p}]$, and $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F_{0}F+f_{0}\|^{2}=\|F\|^{2}+\|f_{0}\|^{2}\}.$